OpenStudy (anonymous):

cos^2x +sinx +1=0 solve for x, please tell me how to do it

OpenStudy (anonymous):

\[\cos^2x +sinx +1=0\]

OpenStudy (anonymous):

i know what the answer is and it is 3pi/2

OpenStudy (anonymous):

but i want to know how to get to it

OpenStudy (anonymous):

cos^2 x + sin x + 1 = 0 1 - sin^2 x + sin x + 1 = 0 - sin^2 x + sin x + 2 = 0 sin^2 x - sin x - 2 = 0

OpenStudy (anonymous):

Then use the substitution u = sin x

OpenStudy (anonymous):

And solve for the quadratic equation, then substitute sin x back into u

OpenStudy (anonymous):

omg thank you!!!